# Is In-Context Learning a Type of Gradient-Based Learning? Evidence from the Inverse Frequency Effect in Structural Priming

### v1

@@ 3.2 Corpus
We adapted the Core Dative Prime-LM Corpus from Sinclair et al. (2022) to create our dataset. We briefly introduce the relevant properties of the corpus and refer the readers to the original paper for details. The dative corpus consists of sentences in two forms:
@@
Each DP is a determiner with a common noun (120 distinct nouns in total). The corpus was constructed in a way that controlled for the degree of semantic association and lexical overlap between prime and target sentences, and sentences are semantically plausible as the ditransitive verbs were manually labeled with their verb frames.
Since our goal is to study the IFE, we will need to ensure variation in the occurrence of verbs with different verb biases in the the prime and target sentences. To do this, for each prime sentence involving one of 22 possible verbs, we sampled 50 target sentences with no lexical overlap. This means that the verbs in prime and target sentences are always distinct. Each prime-target pair can be instantiated structurally in 4 different ways, depending on whether the target and prime sentences are PD or DO: $t_{\text{PD}}|p_{\text{PD}}$, $t_{\text{PD}}|p_{\text{DO}}$, $t_{\text{DO}}|p_{\text{PD}}$, $t_{\text{DO}}|p_{\text{DO}}$ (i.e., target sentence $t$ conditioned on prime $p$). This resulted in 92400 prime-target pairs. [^3] In this paper, we use $t$ and $p$ for individual target and prime sentences, $T$ and $P$ for sets of target and prime sentences, and $\mathcal{P}$ for probability. An example of $t_{\text{PD}}|p_{\text{DO}}$ is “A doctor brought a chief a plate. The secretary drew the card for the band.”
We also created an alternative dataset of the same size by replacing all indirect object DPs with a pronoun. [^4] Details of the set of pronouns and their relative probabilities are in Appendix A. This was motivated by a corpus parse [^5] In order to estimate the verb biases represented in the training corpus of GPT2 models, we parsed a fragment (around 160 million tokens) of the OpenWebText corpus (Gokaslan and Cohen, 2019) with the Python package spaCy (Honnibal et al., 2020) to get an estimate of the DO vs. PD ratio for each verb. We found that the verb biases from the corpus are less well-represented in GPT2 models, motivating our decision to estimate verb biases by using model judgments rather than corpus statistics; see Section 3.4. that showed that the most common indirect objects in DO sentences are animate pronouns, suggesting that it may be important to consider pronouns in analyses of DO vs. PD sentences. The importance of animacy has also previously been noted by Bresnan et al. (2007). The presence and absence of pronouns lead to different verb biases for LLMs, which affect their IFE behaviors. We will return to this point in the discussion.
@@ 3.3 Language Models
We considered a set of Transformer models that have been claimed to show ICL capabilities to various extents (Lee et al., 2024):
  • GPT2 Radford et al. (2019) in three of its sizes (small, medium, large), with 85M, 302M, and 708M parameters, respectively. All versions were loaded from the package transformerLens Nanda and Bloom (2022).
  • LLAMA2 Touvron et al. (2023) in three versions: 7b (6.5B parameters), 7b-chat (6.5B parameters), and 13b (13B parameters). All versions were loaded using the Huggingface transformers library Wolf et al. (2020).
  • GPT3-base Brown et al. (2020) with the davinci-002 version (175B parameters), accessed via the OpenAI API.
The models vary in size, and correspondingly, in their ICL capabilities. We predict that there will be a stronger IFE as size increases. [^6] We also tested the LSTM models from (Gulordava et al., 2018) with the Concatenation mode and found no evidence of structural priming.
Figure 3: A demonstration of the IFE: a stronger priming effect of a DO prime is predicted as PD-bias increases. Note that, because the prime and target sentence differ in structure, the target sentence’s log probability is lower (more negative) when the effect of priming is larger. The numerical values of the primed log probabilities are hypothetical, for illustration purposes only.
@@ 3.4 Quantifying Verb Biases
The verb bias for a specific verb is the likelihood of producing structure $X$ compared to the alternative structure $Y$. In human experiments, baseline verb biases are estimated as the ratio of the number of times one structure occurs over the sum of both structures; such occurrences can be counted either in natural production experiments or in corpus searches Zhou and Frank (2023). Here, we compute a verb bias for each verb analogously as the ratio of the model probability of one structure over the sum of the probabilities of both structures. The probability of a sentence $s$ is the product of probabilities assigned by LLMs to each token $w_{i}$: $\mathcal{P}(s)=\prod_{i}\mathcal{P}(w_{i}|\overline{w}_{<i})$. [^7] In practice, we took the sum of the log probabilities assigned by LLMs to each token in the target sentence, which is equivalent to taking the log of the product of raw probabilities. This measures how likely it is for the model to see or produce this sentence. Then, given a set of sentences $\mathcal{S}_{V}$ with ditransitive verb $V$, where each sentence $t_{\text{PD}}$ with structure PD always has its counterpart $t_{\text{DO}}$ in the opposite structure of DO (see 3.2 and 3.2), the $\mathbf{PD}$-bias of verb $\mathbf{V}$ is the mean normalized probability of sentences in structure $PD$:
$bias(V,PD)=\frac{1}{|\mathcal{S}_{V}|}\sum_{t_{\text{PD}}\in\mathcal{S}_{V}}\frac{\mathcal{P}(t_{\text{PD}})}{\mathcal{P}(t_{\text{PD}})+\mathcal{P}(t_{\text{DO}})}$ (1)
The DO-bias of verb $V,$ i.e., $bias(V,DO)$, is defined analogously. We compute verb bias separately for each model, and use a model’s biases in our assessment of whether it shows the IFE.
@@ 3.5 Simulating Structural Priming
The probability of the target sentence after priming is the product of probabilities assigned to its tokens: $\mathcal{P}(t_{X}|p_{X})=\prod_{i}\mathcal{P}(t_{X_{i}}|p_{X},t_{X_{<i}})$. Under the standard priming effect, the probability of the target sentence $t_{X}$ should be greater after being primed by a ditransitive prime sentence of either structure: $\mathcal{P}(t_{X}|t_{X})>\mathcal{P}(t_{X})$ and $\mathcal{P}(t_{X}|p_{Y})>\mathcal{P}(t_{X})$. Being primed by the same structure induces a larger probability increase than being primed by the opposite structure: $\mathcal{P}(t_{X}|p_{X})>\mathcal{P}(t_{X}|p_{Y})$.
@@ 3.6 Predictions about the Inverse Frequency Effect
Recall that the IFE states that the priming strength of structure $X$ inversely correlates with the prime verb’s $X$-bias. That is, the degree to which the target production’s probability deviates from the baseline is determined by the prime verb. Consider, for example, the case in which the prime sentence has a DO structure and the target sentence has a PD structure (other combinations of structures work analogously). For each prime verb $V$, we computed the PrimeBias for the PD target structure given a DO prime sentence, where PrimeBias is defined as the normalized target probability primed by this verb over a set of target sentences:
@@
As shown in Figure 3, the IFE predicts that with a PD target and DO prime sentence, as the prime verb $V$’s PD-bias increases, the prime sentence is less expected, resulting in a larger priming strength towards the DO direction in the target production, i.e., a smaller $PrimeBias(\text{PD}|\text{DO},V)$ value. Similarly, as PD-bias increases, a PD prime sentence will result in a smaller priming strength towards the PD direction in target production, i.e., again a smaller $PrimeBias(\text{PD}|\text{PD},V)$ value. Therefore, when plotting $PrimeBias(\text{PD}|\text{DO},V)$ and $PrimeBias(\text{PD}|\text{PD},V)$ against increasing verb biases and fitting a line with linear regression, the IFE predicts negative slopes for both plots. Moreover, standard priming predicts that $PrimeBias(\text{PD}|\text{PD},V)$ should have a higher intercept than $PrimeBias(\text{PD}|\text{DO},V)$ since the former increases the probability of $T_{\text{PD}}$ more than the latter. [^8] The other two conditions, namely $T_{\text{DO}}|P_{\text{PD}}$ and $T_{\text{DO}}|P_{\text{DO}}$, are mathematically guaranteed to have slopes that are exactly opposite from $T_{\text{PD}}|P_{\text{PD}}$ and $T_{\text{PD}}|P_{\text{DO}}$, respectively, and the intercept for each should be 1 minus the intercept of its counterpart. We do not plot these other two conditions because these mathematical relationships mean that they provide no information that is not already provided by the plots we do show.
@@ 4 Results and Analysis
Figure 4: GPT2-small shows a robust IFE under the Fine-tuning mode; the negative slope in these plots is the signature of the IFE. Both the $T_{\text{PD}}|P_{\text{PD}}$ (left) and $T_{\text{PD}}|P_{\text{DO}}$ (right) conditions have negative slopes, and the $T_{\text{PD}}|P_{\text{PD}}$ has a higher intercept than $T_{\text{PD}}|P_{\text{DO}}$.
For each model and for each prime verb, we plotted $PrimeBias(\text{PD}|\text{PD},V)$ and $PrimeBias(\text{PD}|\text{DO},V)$ against increasing verb biases and used linear regression to find the relationship between priming strength and verb biases. We report the R-squared ($R^{2}$) coefficient and the root mean squared error (RMSE) to assess the significance of the fitted lines.
@@
Models With Pronoun PDPD_slope PDPD_intercept PDPD_$\mathbf{R^{2}}$ PDPD_RMSE DOPD_slope DOPD_intercept DOPD_$\mathbf{R^{2}}$ DOPD_RMSE GPT2-small True 0.011 0.370 0.014 0.020 -0.007 0.278 0.008 0.017 GPT2-small False 0.014 0.746 0.024 0.016 0.006 0.653 0.003 0.019 GPT2-medium True -0.013 0.351 0.015 0.023 -0.026 0.256 0.107 0.016 GPT2-medium False -0.023 0.748 0.067 0.017 -0.035 0.590 0.060 0.027 GPT2-large True 0.011 0.330 0.017 0.019 -0.037 0.241 0.173 0.018 GPT2-large False -0.003 0.698 0.001 0.018 -0.020 0.487 0.026 0.024 LLAMA2-7b True -0.020 0.392 0.073 0.015 -0.086 0.229 0.645 0.013 LLAMA2-7b False -0.026 0.807 0.046 0.019 -0.111 0.627 0.149 0.042 LLAMA2-7b-chat True -0.012 0.413 0.019 0.018 -0.095 0.263 0.587 0.017 LLAMA2-7b-chat False -0.013 0.788 0.007 0.024 -0.102 0.605 0.107 0.044 LLAMA2-13b True -0.059 0.434 0.323 0.018 -0.099 0.256 0.760 0.011 LLAMA2-13b False -0.066 0.859 0.160 0.019 -0.177 0.685 0.224 0.042 davinci-002 True -0.078 0.403 0.570 0.013 -0.078 0.223 0.662 0.011 davinci-002 False -0.064 0.851 0.172 0.020 -0.145 0.632 0.257 0.035
Table 1: The slope, intercept, $R^{2}$, and RMSE of the fitted lines for each condition under the Concatenation mode. Conditions with both negative slopes are bold (which suggests capturing the IFE), and $R^{2}$ scores higher than $0.5$ are bold (which means a more significant fitted line). PDPD means that the prime and target sentence both have a PD structure, while DOPD means that the prime sentence has a DO structure and the target sentence has a PD structure.
@@ 4.1 Fine-tuning Mode
We applied the Fine-tuning mode to GPT2-small. [^9] We did not carry out this mode for larger models because of the substantial computational resources they require: each priming instance requires a separate fine-tuning process. However, this is unproblematic for our conclusions, given that the Fine-tuning mode is expected to show the IFE in all cases, given its explicit gradient updates. As is shown in Figure 4, the $T_{\text{PD}}|P_{\text{PD}}$ condition has a larger intercept than the $T_{\text{PD}}|P_{\text{DO}}$ condition, suggesting that the Fine-tuning mode is able to capture standard structural priming. Further, we observe negative slopes in both plots, suggesting that the Fine-tuning mode is able to capture the IFE. The $T_{\text{PD}}|P_{\text{DO}}$ condition has a higher $R^{2}$ score of $0.75$, demonstrating a stronger IFE than the $T_{\text{PD}}|P_{\text{DO}}$ condition.
Overall, these results show that even the smallest model we analyzed shows the IFE under explicit gradient-based weight updates, which passes the sanity check by confirming that LLMs are capable of showing the IFE with explicit gradient-based weight updates. We now turn to the crucial case of ICL, to see whether the IFE arises in LLMs even without explicit weight updates.
@@ 4.2 Concatenation Mode
We applied the Concatenation mode to all models. We only show one plot for each of the three types of models (i.e., showing just one variant of that model type and showing only the WithPronoun condition) and report the full results (with and without pronouns across all model variants) in Table 1. As shown in Figure 5, for all models across all conditions, the $T_{\text{PD}}|P_{\text{PD}}$ intercept is greater than the $T_{\text{PD}}|P_{\text{DO}}$ intercept, showing the standard structural priming effect, which is consistent with our prediction. The RMSE score is less than $0.04$ for all conditions, suggesting a significant predictability of the fitted lines to the data points. For the IFE, we found that all three sizes of GPT2 failed to show the IFE, as the slopes are either positive or close to zero. This suggests that, in GPT2, priming strength is not correlated with verb bias under our metric. All three LLAMA2 models showed negative slopes in both cases, which is consistent with the IFE. However, only in the WithPronoun $T_{\text{PD}}|P_{\text{DO}}$ condition are the $R^{2}$ coefficients constantly greater than $0.5$ across the three models, [^10] Given no consensus on standard $R^{2}$ score thresholds, we picked this criterion by default. suggesting that LLAMA2 displays the IFE but in a noisy way that is not fully robust. Finally, for GPT3, both $T_{\text{PD}}|P_{\text{PD}}$ and $T_{\text{PD}}|P_{\text{DO}}$ under the WithPronoun condition have an $R^{2}$ coefficient greater than $0.5$, while $R^{2}$ is not greater than 0.5 in either case in the NoPronoun condition.
Therefore, besides corroborating the finding from prior work that LLMs show structural priming effects, the current results suggest that, in general, larger models tend to show a stronger IFE than smaller models, meaning that IFE strength correlates with ICL capability (since larger models also typically perform better at ICL than smaller models). Assuming that ICL capability correlates with LLM size, given the currently observed pattern, we predict that larger models such as GPT4 should show an even stronger and more significant IFE.
@@
As shown in Table 1, the majority of cases with an $R^{2}$ score above $0.5$ are the WithPronoun $T_{\text{PD}}|P_{\text{DO}}$ cases. Why the IFE arises more strongly in the WithPronoun condition than the NoPronoun condition remains unclear. One major difference between these conditions lies in the default verb biases: as shown in Figure 6, GPT3 shows an overwhelming bias towards PD without pronouns but a reverse pattern favoring DO with pronouns. This pattern holds across all models and is consistent with statistics obtained from our corpus parse. One possible conclusion is that the most common indirect object DPs in the DO sentences are animate pronouns, causing the model to assign a higher probability to pronoun sentences. However, it still remains puzzling why and how these differences in verb biases could lead to the IFE having different levels of significance across the two conditions.
Figure 6: Comparison of PD biases with (left) and without (right) pronouns for GPT3. As is shown in Equation 1, a high PD-bias means a larger proportion of probability mass is assigned to the PD structure than the DO structure in LLMs.

### v2

@@ 3.2 Corpus
We adapted the Core Dative Prime-LM Corpus from Sinclair et al. (2022) to create our dataset. We briefly introduce the relevant properties and refer the readers to the original paper for details. The dative corpus consists of sentences in two forms:
@@
Each DP is a determiner with a common noun (120 distinct nouns in total). The corpus was constructed in a way that ensures semantic plausibility and controls for the degree of semantic association and lexical overlap between prime and target sentences.
Since our goal is to study the IFE, we need to ensure variation in the occurrence of verbs with different verb biases in the the prime and target sentences. To do this, for each prime sentence involving one of 22 possible verbs, we sampled 50 target sentences that have no lexical overlap with the prime sentence. This means that the verbs in prime and target sentences are always distinct. Each prime-target pair can be instantiated structurally in 4 different ways, depending on whether the target and prime sentences are PD or DO: $t_{\text{PD}}|p_{\text{PD}}$, $t_{\text{PD}}|p_{\text{DO}}$, $t_{\text{DO}}|p_{\text{PD}}$, $t_{\text{DO}}|p_{\text{DO}}$ (i.e., target sentence $t$ conditioned on prime $p$). This resulted in 92400 prime-target pairs. [^2] We use $t$ and $p$ for individual target and prime sentences, $T$ and $P$ for sets of targets and primes, and $\mathcal{P}$ for probability. An example of $t_{\text{PD}}|p_{\text{DO}}$ is “A doctor brought a chief a plate. The secretary drew the card for the band.”
We also created an alternative dataset of the same size by replacing each indirect object DP with a pronoun. This was motivated by a corpus study (see Appendix A) that showed that the most common indirect objects in DO sentences are animate pronouns, suggesting that it may be important to consider pronouns in analyses of DO vs. PD sentences. The importance of animacy has also previously been noted by Bresnan et al. (2007). The presence and absence of pronouns lead to different verb biases for LLMs, which affect their IFE behaviors. We return to this point in the discussion.
@@ 3.3 Language Models
We study a set of Transformer models that have been claimed to show ICL capabilities to varying extents (Lee et al., 2024). First, we used GPT2 Radford et al. (2019) in three of its sizes (small, medium, large), with 85M, 302M, and 708M parameters, respectively. All versions were loaded from the package transformerLens Nanda and Bloom (2022). Second, we used Llama2 Touvron et al. (2023) in three versions: 7b (6.5B parameters), 7b-chat (6.5B parameters), and 13b (13B parameters). All versions were loaded using the Huggingface transformers library Wolf et al. (2020). Finally, we used GPT3-base Brown et al. (2020) with the davinci-002 version (175B parameters), accessed via the OpenAI API. The models vary in size, and correspondingly, in their ICL capabilities, as it has been argued that larger models show stronger ICL capability (e.g., Wei et al., 2023; Dong et al., 2024; Chen et al., 2025). We predict that there will be a stronger IFE as size increases. [^3] We also tested the LSTM models from Gulordava et al. (2018) with the Concatenation mode and found no evidence of structural priming.
Figure 2: A demonstration of the IFE: a stronger priming effect of a DO prime is predicted as PD-bias increases. The numerical values of the primed log probabilities are hypothetical, for illustration purposes only.
@@ 3.4 Quantifying Verb Biases
The verb bias for a specific verb is the likelihood of producing structure $X$ compared to the alternative structure $Y$. In human experiments, baseline verb biases are estimated as the ratio of the number of times one structure occurs over the sum of both structures; such occurrences can be counted either in natural production experiments or in corpus searches Zhou and Frank (2023). Here, we compute a verb bias for each verb analogously as the ratio of the model probability of one structure over the sum of the probabilities of both structures. The probability of a sentence $s$ is the product of probabilities assigned by LLMs to each token $w_{i}$: $\mathcal{P}(s)=\prod_{i}\mathcal{P}(w_{i}|\overline{w}_{<i})$.This measures how likely it is for the model to see or produce this sentence. Then, given a set of sentences $\mathcal{S}_{V}$ with ditransitive verb $V$, where each sentence $t_{\text{PD}}$ with structure PD always has its counterpart $t_{\text{DO}}$ in the opposite structure of DO (see 3.2 and 3.2), the $\mathbf{PD}$-bias of verb $\mathbf{V}$ is the mean normalized probability of sentences in structure PD:
$bias(V,\text{PD})=\frac{1}{|\mathcal{S}_{V}|}\sum_{t_{\text{PD}}\in\mathcal{S}_{V}}\frac{\mathcal{P}(t_{\text{PD}})}{\mathcal{P}(t_{\text{PD}})+\mathcal{P}(t_{\text{DO}})}$ (1)
The DO-bias of verb $V,$ i.e., $bias(V,\text{DO})$, is defined analogously. We compute verb bias separately for each model (see Appendix C for GPT3), and use a model’s biases in our assessment of whether it shows the IFE.
@@ 3.5 Simulating Structural Priming
The probability of the target sentence after priming is the product of probabilities assigned to its tokens: $\mathcal{P}(t_{X}|p_{X})=\prod_{i}\mathcal{P}(t_{X_{i}}|p_{X},t_{X_{<i}})$. Under standard priming, the probability of the target sentence $t_{X}$ should be greater after being primed by a ditransitive prime sentence of either structure: $\mathcal{P}(t_{X}|p_{X})>\mathcal{P}(t_{X})$ and $\mathcal{P}(t_{X}|p_{Y})>\mathcal{P}(t_{X})$. Being primed by the same structure should induce a larger probability increase than being primed by the opposite structure: $\mathcal{P}(t_{X}|p_{X})>\mathcal{P}(t_{X}|p_{Y})$.
@@ 3.6 Predictions about the Inverse Frequency Effect
Recall that, under the IFE, the priming strength of structure $X$ inversely correlates with the prime verb’s $X$-bias. That is, the degree to which the target production’s probability deviates from the baseline is determined by the prime verb. Consider, for example, the case in which the prime sentence has a DO structure and the target sentence has a PD structure (other combinations of structures work analogously). For each prime verb $V$, we computed the PrimeBias for the PD target structure given a DO prime sentence, where PrimeBias is defined as the normalized target probability primed by this verb over a set of target sentences:
@@
As shown in Figure 2, the IFE predicts that with a PD target and DO prime sentence, as the prime verb $V$’s PD-bias increases, the prime sentence is less expected, resulting in a larger priming strength towards the DO direction in the target production, i.e., a smaller $PrimeBias(\text{PD}|\text{DO},V)$ value. Similarly, as PD-bias increases, a PD prime sentence will result in a smaller priming strength towards the PD direction in the target production, i.e., again a smaller $PrimeBias(\text{PD}|\text{PD},V)$ value. Therefore, when plotting $PrimeBias(\text{PD}|\text{DO},V)$ and $PrimeBias(\text{PD}|\text{PD},V)$ against increasing PD verb biases and fitting a line with linear regression, the IFE predicts negative slopes for both plots. Moreover, standard priming predicts that $PrimeBias(\text{PD}|\text{PD},V)$ should have a higher intercept than $PrimeBias(\text{PD}|\text{DO},V)$ since the former increases $\mathcal{P}(T_{\text{PD}})$ more than the latter. [^4] The other two conditions, namely $T_{\text{DO}}|P_{\text{PD}}$ and $T_{\text{DO}}|P_{\text{DO}}$, are mathematically guaranteed to have slopes that are exactly opposite from $T_{\text{PD}}|P_{\text{PD}}$ and $T_{\text{PD}}|P_{\text{DO}}$, respectively, and the intercept should be 1 minus the intercept of its counterpart.
@@ 4 Results and Analysis
For each model and for each prime verb, we plotted $PrimeBias(\text{PD}|\text{PD},V)$ and $PrimeBias(\text{PD}|\text{DO},V)$ against increasing verb biases and used linear regression to find the relationship between priming strength and verb biases. We report the $95\%$ confidence intervals of the fitted slopes to assess the significance of the fitted lines, where intervals below 00 suggest that the slopes are more likely to be negative.
Figure 3: GPT2-small shows a robust IFE under the Fine-Tuning mode, shown by the two negative slope.
@@ 4.1 Fine-Tuning Mode
We applied the Fine-Tuning mode to GPT2-small. As is shown in Figure 3, the $T_{\text{PD}}|P_{\text{PD}}$ condition has a larger intercept than the $T_{\text{PD}}|P_{\text{DO}}$ condition, suggesting that the Fine-Tuning mode is able to capture standard structural priming. Further, we observe negative slopes in both lines, suggesting that the Fine-Tuning mode is able to capture the IFE. The $95\%$ confidence interval for the $T_{\text{PD}}|P_{\text{DO}}$ condition is below 00, demonstrating a stronger IFE than the $T_{\text{PD}}|P_{\text{PD}}$ condition.
Overall, these results show that even the smallest model we analyzed shows the IFE under explicit error-driven weight updates. We did not carry out this mode for larger models as fine-tuning on the demonstration with GPT2-small already satisfies its sanity-check purpose: as already noted, we have strong reason to expect that the Fine-Tuning mode will show the IFE, given its explicit gradient-based updates. We now turn to the crucial case of ICL, to see whether the IFE arises even without explicit weight updates.
@@ 4.2 Concatenation Mode
We applied the Concatenation mode to all models. In Figure 4, we show one plot for each of the three types of models (i.e., showing just one variant of that model type and showing only the WithPronoun condition) and report the results across all model variants in Table 1 (see Appendix D for the full statistical results). For all models across all conditions, the $T_{\text{PD}}|P_{\text{PD}}$ intercept is greater than the $T_{\text{PD}}|P_{\text{DO}}$ intercept, showing the standard structural priming effect, which is consistent with our prediction. For the IFE, we found that all three sizes of GPT2 failed to show the IFE, as the slopes are either positive or close to zero. This suggests that, in GPT2, priming strength is not correlated with verb bias under our metric. All three Llama2 models showed negative slopes in both cases, which is consistent with the IFE. However, only in the WithPronoun $T_{\text{PD}}|P_{\text{DO}}$ condition are the $95\%$ confidence intervals constantly below 00 across the three models, suggesting that Llama2 displays the IFE but in a noisy way that is not fully robust. Finally, for GPT3, both $T_{\text{PD}}|P_{\text{PD}}$ and $T_{\text{PD}}|P_{\text{DO}}$ under the WithPronoun condition have their confidence intervals below 00, while the confidence intervals in both case in the NoPronoun condition contain 00.
The WithPronoun DO-PD conditions of all three Llama2 models and GPT3 are statistically significant with $p<0.0001$. Models at least as large as Llama2-13b show at least marginal statistical significance in all conditions while smaller models do not. These results support our hypothesis that larger models show a more significant IFE.
Therefore, besides corroborating the finding from prior work that LLMs show structural priming effects, the current results suggest that larger models tend to show a stronger IFE than smaller models, meaning that IFE strength correlates with ICL capability since larger models also typically perform better at ICL than smaller models (Brown et al., 2020). Assuming that ICL capability correlates with LLM size, given the currently observed pattern, we predict that larger models such as GPT4 should show an even stronger IFE.
Intercept Slope $95\%$ CI of Slopes Model Pronoun Obj? PD-PD DO-PD PD-PD DO-PD PD-PD DO-PD GPT2-small True 0.370 0.278 0.011 -0.007 (-0.049, 0.072) (-0.058, 0.044) GPT2-small False 0.746 0.653 0.014 0.006 (-0.042, 0.07) (-0.06, 0.072) GPT2-medium True 0.351 0.256 -0.013 -0.026 (-0.078, 0.053) (-0.073, 0.022) GPT2-medium False 0.748 0.590 -0.023 -0.035 (-0.079, 0.032) (-0.125, 0.054) GPT2-large True 0.330 0.241 0.011 -0.037+ (-0.043, 0.065) (-0.09, 0.015) GPT2-large False 0.698 0.487 -0.003 -0.02 (-0.062, 0.055) (-0.098, 0.058) Llama-7b True 0.392 0.229 -0.02 -0.086∗∗∗∗ (-0.067, 0.026) (-0.126, -0.045) Llama-7b False 0.807 0.627 -0.026 -0.111+ (-0.102, 0.049) (-0.279, 0.057) Llama-7b-chat True 0.413 0.263 -0.012 -0.095∗∗∗∗ (-0.067, 0.043) (-0.146, -0.044) Llama-7b-chat False 0.788 0.605 -0.013 -0.102 (-0.115, 0.089) (-0.289, 0.085) Llama-13b True 0.434 0.256 -0.059∗∗ -0.099∗∗∗∗ (-0.113, -0.005) (-0.134, -0.063) Llama-13b False 0.859 0.685 -0.066+ -0.177 (-0.163, 0.031) (-0.388, 0.033) GPT3-davinci-002 True 0.403 0.223 -0.078∗∗∗∗ -0.078∗∗∗∗ (-0.121, -0.035) (-0.114, -0.043) GPT3-davinci-002 False 0.851 0.632 -0.064+ -0.145 (-0.153, 0.025) (-0.301, 0.012)
Table 1: The intercept, slope, and $95\%$ confidence interval of the fitted lines for each condition under the Concatenation mode. We mark the $p$-values of the fitted slopes with the following notation: * for $p\leq 0.05$, ** for $p\leq 0.01$, **** for $p\leq 0.0001$, and $+$ for $p\leq 0.1$ (marginally significant). $95\%$ confidence intervals of the fitted slopes below zero are bold (which means the fitted slopes are more statistically significant to be negative).
@@
As shown in Table 1, the majority of cases with their $95\%$ confidence intervals excluding 00 are the WithPronoun $T_{\text{PD}}|P_{\text{DO}}$ cases. Why the IFE arises more strongly in the WithPronoun condition than the NoPronoun condition remains unclear. One major difference between these conditions lies in the default verb biases: as shown in Figure 5 in Appendix C, GPT3 shows an overwhelming bias towards PD without pronouns but a reverse pattern favoring DO with pronouns. This pattern holds across all models and is consistent with statistics obtained from our corpus parse. One possible conclusion is that the most common indirect object DPs in the DO sentences are animate pronouns, causing the model to assign a higher probability to pronoun sentences. The asymmetry between the WithPronoun and NoPronoun conditions can be explained by our claim that ICL becomes more like error-driven learning with both model scale and data scale (see Section 3.1). The role of model scale has already been demonstrated in Table 1 (i.e., larger models show stronger IFE in general); here we elaborate on the role of data scale, which is somewhat more subtle.
In the corpus study mentioned in Appendix A, we found that pronominal indirect object (IO) sentences are more common than non-pronominal ones in the OpenWebText corpus. Since larger models have seen more pronominal IO sentences compared to non-pronominal ones, they are expected to perform more like error-driven learning for pronominal IO sentences than non-pronominal IO sentences in ICL, under the assumption that data scale (like model scale) increases the strength of the IFE. Such relationships between data scale and the strength of ICL-related effects have also been observed in McCoy et al. (2024), who showed that LLMs displayed better ICL performance on high-probability sentences than low-probability ones. This observation could also potentially explain the interesting finding from Sinclair et al. (2022) that structural priming in LLMs (to a greater degree than in humans) is modulated by semantic plausibility: semantically plausible inputs are better represented in the training data, leading to more ICL/structural priming.
One further empirical observation from our experiments supports this reasoning: for Llama2-13B and GPT3, we observed (marginally) significant IFE (p < 0.1 or p < 0.05) for the NoPronoun conditions, although it is a smaller effect than in the WithPronoun conditions. The fact that there are significant effects in both cases suggests that the asymmetry between the two conditions is quantitative rather than qualitative. This graded view aligns well with the data-scale-based account: in addition to the difference of IFE significance explained by model scale, the across-condition differences within single models are accounted for by the fact that the model is trained on more with-pronoun dative sentences than without-pronoun ones, which makes the model better approximate error-driven learning in ICL with the with-pronoun ones, which leads to a more significant IFE as we have observed.